# Predicting random close packing of binary hard-disk mixtures via third-virial-based parameters

### v1

@@ Acknowledgements.
A.S acknowledges financial support from Grant No. PID2024-156352NB-I00 funded by MCIU/AEI/10.13039/501100011033/FEDER, UE and from Grant No. GR24022 funded by the Junta de Extremadura (Spain) and by European Regional Development Fund (ERDF) “A way of making Europe.”
@@
Appendix
Here we show that Eqs. (1) are the only possible forms if one assumes that $1/\phi_{\text{mixt}}$ is a linear function of $\mu_{B}$ and that $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ is a linear function of $\lambda_{B}\equiv 1/(1-\mu_{B})$. Under these assumptions one may write
@@
where we have used that $\mu_{B}=0$ in the monocomponent limit. Combining Eqs. (10) one obtains
@@
Since this identity must hold for arbitrary $\mu_{B}$, both the constant and linear terms must vanish independently. This yields $A=1/B=(1-\phi_{\text{mono}})/\phi_{\text{mono}}$, in agreement with Eqs. (1).
@@
Eliminating $\phi_{\text{mixt}}$ between Eqs. (12) gives
@@
Requiring this expression to be valid for all $\mu$ leads to $C=1/D=1-\phi_{\text{mono}}$, in agreement with Eqs. (2).

### v3

@@ Acknowledgements.
A.S. acknowledges financial support from Grant No. PID2024-156352NB-I00 funded by MCIU/AEI/10.13039/501100011033 and by ERDF/EU, and from Grant No. GR24022 funded by the Junta de Extremadura (Spain).
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Andrés Santos: Conceptualization (lead); Formal analysis (equal); Funding acquisition (lead); Investigation (equal); Methodology (equal); Supervision (lead); Writing – original draft (equal); Writing – review & editing (equal). Mariano López de Haro: Formal analysis (equal); Investigation (equal); Methodology (equal); Writing – original draft (equal); Writing – review & editing (equal).
@@
Appendix A Linear dependencies of $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$
Here we show that Eq. (1) represents the only possible forms if one assumes that $1/\phi_{\text{mixt}}$ is a linear function of $\mu_{B}$ and that $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ is a linear function of $\lambda_{B}\equiv 1/(1-\mu_{B})$. Under these assumptions, one may write
@@ Appendix A Linear dependencies of $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$
where we have used the fact that $\mu_{B}=0$ in the monocomponent limit. Combining Eq. (12) one obtains
@@
Since this identity must hold for arbitrary $\mu_{B}$, both the constant and linear terms must vanish independently. This yields $A=1/B=(1-\phi_{\text{mono}})/\phi_{\text{mono}}$, in agreement with Eq. (1).
@@
Eliminating $\phi_{\text{mixt}}$ between Eq. (14) gives
@@
Requiring this expression to be valid for all $\mu$ leads to $C=1/D=1-\phi_{\text{mono}}$, in agreement with Eq. (2).
Appendix B The surplus equation of state for mixtures
By imposing a set of consistency conditions, a fundamental-measure-theory approach linking the equilibrium equation of state of a three-dimensional monocomponent HS fluid to that of a three-dimensional polydisperse HS mixture was derived.[94, 95, 48, 79] In this so-called surplus approximation, the compressibility factor of the mixture, $Z_{\text{mixt}}$, is expressed in terms of that of the monocomponent fluid, $Z_{\text{mono}}$, as
$Z_{\text{mixt}}(\phi_{\text{mixt}})=\alpha\frac{\phi_{\text{mono}}}{\phi_{\text{mixt}}}Z_{\text{mono}}(\phi_{\text{mono}})+\frac{1-\alpha/\lambda}{1-\phi_{\text{mixt}}},$ (16)
where the packing fraction $\phi_{\text{mono}}$ of the monocomponent fluid is related to the packing fraction $\phi_{\text{mixt}}$ of the polydisperse mixture through
$\phi_{\text{mono}}=\frac{\phi_{\text{mixt}}}{\phi_{\text{mixt}}+\lambda(1-\phi_{\text{mixt}})},$ (17a)
$\frac{\phi_{\text{mixt}}}{1-\phi_{\text{mixt}}}=\lambda\frac{\phi_{\text{mono}}}{1-\phi_{\text{mono}}}.$ (17b)
The parameters $\lambda$ and $\alpha$ can be determined by imposing consistency with the (reduced) second and third virial coefficients of the mixture,[79] leading to
$\lambda=\frac{\bar{B}_{2}-1}{b_{2}-1}\frac{b_{3}-2b_{2}+1}{\bar{B}_{3}-2\bar{B}_{2}+1},\quad\alpha=\lambda^{2}\frac{\bar{B}_{2}-1}{b_{2}-1}.$ (18)
Although initially derived for three-dimensional HS systems,[95, 48] Eqs. (16)–(18) can be generalized to any dimensionality $d\neq 3$.[80, 81] In the particular case of HD systems ($d=2$), one has $b_{2}=2$ and $\bar{B}_{2}=1+m_{2}^{-1}$, so that the parameter $\lambda$ reduces to the expression given in Eq. (2).
While the surplus equation of state is restricted to equilibrium states, we borrow two structural features from Eq. (17b) in our proposal to relate the RCP of the HD mixture ($\phi_{\text{mixt}}$) to that of the monocomponent system ($\phi_{\text{mono}}$). First, we assume that all the details of the mixture composition are entirely encapsulated in the universality parameter $\lambda$. Second, we assume that the ratio $\phi_{\text{mixt}}/(1-\phi_{\text{mixt}})$ is a linear function of $\lambda$. To these two assumptions, we add the condition that $\phi_{\text{mixt}}$ is a linear function of $\mu=1-\lambda^{-1}$. As shown in Appendix A, the combination of these three conditions uniquely leads to Eq. (2).